Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 172     Entry time: 21 Sep 2024, 09:50
Author: Peter Zalden 
Group: FXE 
Subject: SA1 status 

The xray beam drifted significantly on wednesday and thursday (18. and 19.) but then was not drifting on friday (20.), see attached trendline of the M2 actuator that kept the beam at fixed position on M3.

Today also seems that beam pointing is stable. But unclear what happened earlier...

Attachment 1: 2024-09-21-094858_073864177_exflqr51474.png  81 kB  Uploaded 21 Sep 2024, 10:51  | Hide | Hide all
2024-09-21-094858_073864177_exflqr51474.png
ELOG V3.1.4-7c3fd00